OpenStudy (anonymous):

Given the sequence 5, -10, 20, -40, the next two numbers in the sequence are 80 and -160. true or false?

OpenStudy (raden):

true

OpenStudy (whpalmer4):

Divide each number in the sequence by the one before it. -10/5 = -2. 20/-10 = -2. -40/20 = -2. See the pattern? Going forward, multiply the last number by -2 to get the next one in the sequence. -40 * (-2) = ?

OpenStudy (whpalmer4):

Because there is a constant ratio between successive terms in the sequence (in this case, -2) this is a geometric sequence or progression, and that constant ratio is called the common ratio.

OpenStudy (anonymous):

\[\left\{ 5n*(-1)^{n+1} \right\} \] is the equation for the sequence, plugging in the correct term at n, would provide you insight, the sequence starts at t = 1

OpenStudy (anonymous):

I meant n = 1! not t. Sorry